# thanks: All the authors contributed equallythanks: All the authors contributed equallythanks: All the authors contributed equallythanks: All the authors contributed equally

### v2

@@ Abstract
The study of high-dimensional dynamical systems is essential for advancing our understanding of various complex phenomena [e.g., 1, 2, 3, 4, 5, 6, 7, 8]. These systems are characterized by numerous interacting degrees of freedom, manifesting feedback mechanisms across spatial and temporal scales. Notable examples of such complexity are observed in climate modeling [e.g., 9], where feedback mechanisms lead to self-sustained spatio-temporal patterns, such as the El Niño Southern Oscillation (ENSO) in the tropical Pacific Ocean, the Asian Monsoon (particularly prominent in South and East Asia), the Indian Ocean Dipole, and the Madden-Julian Oscillation in the Indian and Pacific Oceans [10, 11], among others. Similar complexities and the need to understand feedback mechanisms can be observed in other areas, such as financial markets, [e.g., 12] or neuroscience [e.g., 13, 14].
@@
A central challenge is to characterize the causal relationships among these degrees of freedom without prior knowledge of the underlying evolution laws [15, 16, 17, 18]. Causal inference seeks to unambiguously determine whether the behavior of one variable has been influenced by another based on observed time series data. This estimation process is challenging due to the complexity and high dimensionality of these systems [19, 20].
@@
In recent decades, numerous methodologies have been developed to infer causal relationships directly from data [e.g., 21, 22, 23, 24, 25], but calculating a system’s response to small perturbations using linear response theory has become a dominant approach [26, 27, 28]. This method allows for evaluating responses without actually perturbing the system, leveraging instead the analysis of unperturbed dynamics [29, 30]. It is underpinned by the generalized Fluctuation-Dissipation Theorem (GFDT), an extension of the classical Fluctuation-Dissipation Theorem (FDT) [31]. A central obstacle in applying linear response theory is obtaining the system’s unperturbed invariant distribution, i.e., the distribution of a system’s attractor [32, 33, 34].
@@
The invariant distribution is often approximated by assuming a multivariate Gaussian distribution ansatz to circumvent this issue [35]. This approximation is reasonable when examining coarse-grained (spatially or temporally) variables or particular physical observables [36, 37]. The limits of this approximation are evident in atmospheric science, where observational data from wind and rainfall exhibit intermittency and other non-Gaussian features [38, 39, 40, 41]. Another context for non-Gaussian statistics comes from neuroscience, where the brain’s response to stimuli and the resulting patterns of neural activity display non-Gaussian properties [42].
@@
Our work leverages recent advancements in score-based generative modeling [e.g., 43, 44, 45, 46, 47, and follow-on studies] to accurately represent a high-dimensional, nonlinear dynamical system’s invariant distribution via a “score function,” the gradient of the logarithm of a system’s invariant distribution. Score models are used primarily as generative models, i.e., for sampling from the data distribution. Still, they have other applications, such as estimating the dimensionality of the data distribution, see [48]. This study demonstrates another use case beyond sample generation and in an area where machine-learning approaches have not yet been fully utilized: a trained score model can be used to compute response functions using data from dynamical systems.
@@
Using trajectories of the dynamical system, we train a model that approximates the score of the steady-state, or invariant, distribution. The model is a neural network based on spatial convolutions that can capture high-dimensional data distribution statistics (e.g., [49, 50, 51]). Like the Gaussian approximation approach, this method only requires time-series data. In this work, we apply the score-based method to three dynamical systems and compare its performance to other standard approaches for computing response functions. This approach outperforms the traditional Gaussian approximation when the underlying system is nonlinear.
@@
For a small perturbation to $u$ at grid location $j$, denoted by $\delta u_{j}$, we wish to quantify the change in the average of $u_{i}$ at a later time $\mathcal{T}$ (the discrete analog of Equation (2)). The GFDT states that this can be evaluated using a matrix-valued quantity called the response function $R_{ij}$ (we also refer to it as the response matrix) [31] as
@@
Since we assume that we have access to time-series data, the problem of estimating the response function is thus reduced to estimating the score function $\vec{s}(\vec{u})$ of the steady-state distribution $\rho$. In deterministic dynamical systems, this invariant distribution is typically singular almost everywhere on the attractor, and it is necessary to introduce Gaussian noise into the system to make $\rho$ sufficiently smooth to apply the GFDT [52] effectively.
@@
where $\bm{C}(\mathcal{T})$ is the correlation matrix, with elements $C_{ij}(\mathcal{T})=\langle u_{i}(\mathcal{T})u_{j}(0)\rangle_{\mathcal{J}}$. The property that $R_{ij}(0)=\delta_{ij}$, where $\delta_{ij}$ is the Kronecker delta, is a general property that holds for the discrete response matrix, see the Appendix. Constructing the correlation matrix is straightforward, allowing Equation (5) to serve as a convenient method for estimating the response matrix when the underlying steady-state distribution is well-approximated by a Gaussian distribution.
@@
Here, we employ score-based generative modeling, a technique from the machine-learning community, to derive the response function directly from data. Our approach uses denoising score matching and facilitates the parameterization of the dynamical system’s attractor score function directly, following methodologies suggested by Song et al. [45, 46]. This approach only requires data and notably does not require approximating the steady-state distribution as Gaussian or require additional numerical simulations for different perturbations.
Score-based modeling for response functions – In most applications of generative modeling, the goal is to generate samples from the data distribution $\rho$. Generating samples is achieved using a reverse diffusion process, which turns samples from a normal distribution into samples of the data-distribution. This sample generation procedure requires knowledge of the score function from a non-stationary stochastic forward process at each moment in time; one example of such a forward process is
$\frac{\partial\rho_{\tau}}{\partial\tau}=\frac{\sigma^{2}(\tau)}{2}\Delta\rho_{\tau},$ (6)
where $\tau$ is a pseudo-time, and $\sigma(\tau)$ is a prescribed function [46].
The learning objective is to fit the parameters of a model (often a neural network) to approximate $\vec{s}_{\tau}=\nabla\ln\rho_{\tau}$, which appears in the reverse diffusion equation. Knowledge of this function at each diffusion time $\tau$ allows one to sample from a Gaussian, evolve the reverse dynamical system, and obtain a sample from the data distribution. What is relevant for response functions is that taking off-the-shelf score-modeling infrastructure allows one to learn $\vec{s}_{\tau}$, and evaluating this function as $\tau\rightarrow 0$ yields the score of the data distribution (Equation (4)).
Figure 1: Example snapshots of the Three Systems. In the top row, Ornstein-Uhlenbeck PDE, in the middle row Allen-Cahn, in the bottom row Navier-Stokes
We apply the proposed method to three different dynamical systems. We compare the response function computed in this way (Generative) to a response computed by solving the perturbed dynamical system (Dynamical), a response using an analytically derived score function (Analytical) for the systems for which it is possible to derive one, and the response obtained through the Gaussian approximation frequently used in the literature (Gaussian).
@@
An Ornstein-Uhlenbeck Stochastic PDE. – Our first system is an Ornstein-Uhlenbeck (OU) system [53] with spatially correlated noise, [54, 55, 56, 57]. The steady-state trajectories of the system exhibit Gaussian statistics, which makes the Gaussian approximation exact, and hence serves as a baseline for testing our method.
@@
A modified Allen-Cahn system. – Our second system is the Allen-Cahn system, an example of a reaction-diffusion equation used for modeling phase separation in multi-component systems [see 58, who studied applications to binary alloys]. Reaction-diffusion equations [59] are partial differential equations combining a diffusive term with a (non)linear reaction term. They are commonly used to study pattern formation (e.g., [60, 61, 62]) and chemical processes.
@@
For both the OU and Allen-Cahn systems, the analytic score function and the machine-learned score will differ by a constant factor due to the difference between the Kronecker delta and the Dirac delta functions. In both examples, the analytic score function is $32^{2}$ larger. We divide the analytic score by this factor for comparison with the machine-learned score. Refer to the Appendix for the derivation of the analytic score functions.
Navier-Stokes equation. – We next study a version of the 2D Navier-Stokes model as described in [63], where we take the state variable to be the vorticity, $u$. This system was chosen as a stepping-stone towards the more complex fluids observed in climate systems; however, the vorticity statistics are not strongly non-Gaussian due to the chosen resolution as would be expected from a coarse-grained system, [36]. Hence, we expect good performance from the Gaussian approximation to the response function. The data-generating equation is
@@
$\mathcal{D}u$ $=-\nu_{h}\Delta^{-2}u-\nu\Delta^{2}u-\nu_{0}\int\text{d}x\text{d}yu$ (14)
@@
where $\mathcal{D}$ is a dissipation operator with constants ($\nu_{h}=10^{-2}$, $\nu=10^{-5}$, $\nu_{0}=(2\pi)^{-2}$), $A=\pi$ is a mean advection term, and $\varsigma$ is a random wave forcing from [63]. The random wave forcing introduces hidden dynamics into the system, thus the score function is estimated using only partial state information. The inverse Laplacian is defined to preserve the mean of the original variable.
Figure 2: Response functions for Three Systems. The response to a perturbation at the pixel component $i=1$ is evaluated at various pixel components with coordinates indicated at each panel. All evaluated components align with the advection direction. Responses are computed by integrating the dynamical system (orange lines, “Dynamical”) via Gaussian approximation (black lines, “Gaussian”), using the score function derived from generative modeling (red lines and dots, “Generative”), and using the analytic score function (blue lines, “Analytic”).
Figure 3: Response Function Errors for the Three Systems. Here, we show both the RMS error and the Infinity Norm error of the Response function as a function of time. Here we use the Dynamic Response function as the “ground truth”. We see that the Generative response (red) generally outperforms the Gaussian response (blue) for the nonlinear cases (Allen-Cahn and Navier-Stokes), especially in the Infinity Norm, and performs similarly for the linear system (Ornstein-Uhlenbeck). The response function computed with the exact score does not have zero error relative to the dynamical response function because of errors introduced in numerical discretization and empirical averaging.
@@
Results. – We first generate time series data corresponding to Equations (7),  (10), and Equation (12). These data are used to compute response functions and the training sets for the score model. To create the training data, we determine autocorrelation times for each and subsample the trajectories to create statistically independent snapshots. Figure (9) shows training data samples of the numerically simulated systems. All details regarding the numerical simulations (time-stepping, spatial discretization), the datasets and the training of the score models are provided in the Appendix, along with an in-depth comparison between data samples generated via numerical simulation of the Allen-Cahn system and those produced using the learned score function of this system through reverse diffusion. Training the neural network took a few hours to a few days to for each score model on a single Nvidia A100.
@@
Owing to the periodic boundary conditions and symmetries of the systems under study, the response matrix elements exhibit translational invariance in the $x$ and $y$-directions. However, due to the advection term, the response function in the $x$-direction depends on both the distance and the direction (sign) between pixels while the $y$-direction only depends on the distance. This invariance implies that the index $i$ in the response matrix $R_{ij}$ can be arbitrarily fixed. We chose $j=1$ and calculated the response matrix elements along the direction of advection (denoted by $i$), as this is the direction information propagation, where we observe a more significant response to perturbations at pixel component $i=1$. The resulting responses of nearby pixels as a function of time for the three systems are presented in Figure (4). The generative model is consistent with the dynamical response function (“our ground truth”) in all cases. For the OU system, the Gaussian response function also matches the ground truth; this is expected as this system exhibits Gaussian statistics. For the Allen-Cahn system, it is clear that the Gaussian approximation is a poor approximation to the response function. And lastly, for the Navier-Stokes system we see improvement in using the generative score over the Gaussian response, despite the coarse Navier-Stokes system being quasi-Gaussian.
@@
In Figure (5), we compared the error over the entire state field between the Generative, Gaussian, and Analytic (for the systems where it is available) response functions with the Dynamical response function considered as “ground truth” for the three systems as a function of time. We used two different metrics to determine the errors: the $L^{2}$ and the $L^{\infty}$ norm. The $L^{2}$ norm measures the average error across all pixels in the domain, providing a sense of the overall error distribution. It is helpful in understanding the general accuracy of the response function over the entire field. On the other hand, the $L^{\infty}$ norm measures the maximum error at any single pixel, highlighting the worst-case error. This norm is particularly relevant to ensure the minimization of the largest deviations. From the comparison in Figure (5), it is evident that the Generative response function generally outperforms the Gaussian response in the nonlinear cases (Allen-Cahn and Navier-Stokes), especially in terms of the $L^{\infty}$ norm at early to intermediate times. This observation indicates that the Generative model is better at capturing the extreme responses at individual pixels. The Gaussian approximation is exact for the Ornstein-Uhlenbeck system, which has a Gaussian steady-state distribution, and both the Generative and Gaussian response functions perform similarly.
@@
Drawbacks and Practical Considerations. – While score-based generative models offer a powerful tool for estimating response functions, there are practical challenges. One issue is that training these models requires substantial data or a sufficiently good inductive bias in the parameterized model for the score function. Additionally, setting up and training neural networks can often require significant resources, engineering, and time. Expertise in training neural networks is essential. For example, suppose the data or training time is insufficient, and one is using a U-Net to represent the score function. In that case, the generative response function tends to match the Gaussian approximation’s performance due to the U-Net’s linear bypass connections. A further drawback with any machine-learning method is the lack of performance “out-of-sample”; indeed, in the Appendix, we demonstrate that the score model for the Allen-Cahn system performs poorly when evaluated on data dissimilar from the training data.
@@
For many applications, especially those mentioned in the introduction, several additional steps may need to be taken to apply the methods from this study effectively. One crucial step is the incorporation of domain-specific knowledge into the design of the neural network architecture. This will significantly improve the model’s inductive bias and alleviate the data burden. Additionally, domain-specific challenges such as non-stationarity or incomplete state information must be incorporated [64]. This is particularly salient for climate data, which exhibits diurnal, seasonal, and longer timescale variations [65, 66]. As always, it is crucial to validate the trained model with known benchmarks or analytical solutions.
@@
Conclusions. – In this study, we use score-based diffusion models to compute response functions, offering a purely data-driven method to calculate accurate response functions. We validated this approach on a system with Gaussian statistics (Ornstein-Uhlenbeck) and two dynamical systems with more complex distributions: a reaction-diffusion equation (Allen-Cahn) and the Navier-Stokes equations. The former is a class of SPDEs that have applications across biology, chemistry, and physics, and the latter is a system exhibiting some of the characteristics of more complex fluids like the Earth’s atmosphere or ocean. We compared the response function of these systems computed in multiple ways: using an analytic expression for the score, using the generative score function, using a dynamical approach by introducing perturbations to initial conditions and solving the equations of motion, and using a Gaussian approximation. In this case, we found that the Gaussian response function over-predicted the response to neighboring perturbations and that the generative response function agreed well with both the dynamical and analytical response functions. Compared with the Gaussian response, the RMSE of the generative response function had significantly reduced error. This suggests that score-based models are a useful tool for studying system response functions.

### v3

@@ Abstract
The study of high-dimensional dynamical systems is essential for advancing our understanding of various complex phenomena (e.g., Mucha et al., 2010; Halu et al., 2013; Jumper et al., 2021; Brunton et al., 2019; Barbier, 2020; Singh et al., 2021; Dorogovtsev et al., 2008; Barabási and Albert, 1999). These systems are characterized by numerous interacting degrees of freedom, manifesting feedback mechanisms across spatial and temporal scales. Notable examples of such complexity are observed in climate modeling (e.g., Ghil and Lucarini, 2020), where feedback mechanisms lead to self-sustained spatio-temporal patterns, such as the El Niño Southern Oscillation (ENSO) in the tropical Pacific Ocean, the Asian Monsoon (particularly prominent in South and East Asia), the Indian Ocean Dipole, and the Madden-Julian Oscillation in the Indian and Pacific Oceans (Timmermann et al., 2018; Martin et al., 2021), among others. Similar complexities and the need to understand feedback mechanisms can be observed in other areas, such as financial markets, (e.g., Badwan, 2022) or neuroscience (e.g., Hunt et al., 2021; Sporns and Betzel, 2021).
@@
A central challenge is to characterize the causal relationships among these degrees of freedom without prior knowledge of the underlying evolution laws (BozorgMagham et al., 2015; Lagemann et al., 2023; Wismüller et al., 2021; Keyes et al., 2023). Causal inference seeks to unambiguously determine whether the behavior of one variable has been influenced by another based on observed time series data. This estimation process is challenging due to the complexity and high dimensionality of these systems (Aurell and Del Ferraro, 2016; Friedrich et al., 2011).
@@
In recent decades, numerous methodologies have been developed to infer causal relationships directly from data (e.g., Granger, 1969; Schreiber, 2000; Pearl, 2009; Camps-Valls et al., 2023; Kaddour et al., 2022), but calculating a system’s response to small perturbations using linear response theory has become a dominant approach Baldovin et al. (2022); Cecconi et al. (2023); Falasca et al. (2024). This method allows for evaluating responses without actually perturbing the system, leveraging instead the analysis of unperturbed dynamics (Baldovin et al., 2020; Lucarini, 2018). It is underpinned by the generalized Fluctuation-Dissipation Theorem (GFDT), an extension of the classical Fluctuation-Dissipation Theorem (FDT) Marconi et al. (2008). A central obstacle in applying linear response theory is obtaining the system’s unperturbed invariant distribution, i.e., the distribution of a system’s attractor (Daum, 2005; Sjöberg et al., 2009; Cooper and Haynes, 2011).
@@
The invariant distribution is often approximated by assuming a multivariate Gaussian distribution ansatz to circumvent this issue (Leith, 1975). This approximation is reasonable when examining coarse-grained (spatially or temporally) variables or particular physical observables (Sardeshmukh and Sura, 2009; Gritsun and Lucarini, 2017). The limits of this approximation are evident in atmospheric science, where observational data from wind and rainfall exhibit intermittency and other non-Gaussian features (Proistosescu et al., 2016; Loikith and Neelin, 2015; Branicki and Majda, 2012; Martinez-Villalobos and Neelin, 2019). Another context for non-Gaussian statistics comes from neuroscience, where the brain’s response to stimuli and the resulting patterns of neural activity display non-Gaussian properties (Kriegeskorte and Wei, 2021).
@@
Our work leverages recent advancements in score-based generative modeling (e.g., Vincent, 2011; Ho et al., 2020; Song and Ermon, 2019; Song et al., 2020; Bischoff and Deck, 2023; Liu et al., 2023, and follow-on studies) to accurately represent a high-dimensional, nonlinear dynamical system’s invariant distribution via a “score function,” the gradient of the logarithm of a system’s invariant distribution. Score models are used primarily as generative models, i.e., for sampling from the data distribution. Still, they have other applications, such as estimating the dimensionality of the data distribution, see (Stanczuk et al., 2023). This study demonstrates another use case beyond sample generation and in an area where machine-learning approaches have not yet been fully utilized: a trained score model can be used to compute response functions using data from dynamical systems.
@@
Using trajectories of the dynamical system, we train a model that approximates the score of the steady-state, or invariant, distribution. The model is a neural network based on spatial convolutions that can capture high-dimensional data distribution statistics (e.g., (Ronneberger et al., 2015; Adcock et al., 2020; Beneventano et al., 2021)). Like the Gaussian approximation approach, this method only requires time-series data. In this work, we apply the score-based method to three dynamical systems and compare its performance to other standard approaches for computing response functions. This approach outperforms the traditional Gaussian approximation when the underlying system is nonlinear.
@@
For a small perturbation to $u$ at grid location $j$, denoted by $\delta u_{j}$, we wish to quantify the change in the average of $u_{i}$ at a later time $\mathcal{T}$ (the discrete analog of Equation (2)). The GFDT states that this can be evaluated using a matrix-valued quantity called the response function $R_{ij}$ (we also refer to it as the response matrix) Marconi et al. (2008) as
@@
Since we assume that we have access to time-series data, the problem of estimating the response function is thus reduced to estimating the score function $\vec{s}(\vec{u})$ of the steady-state distribution $\rho$. In deterministic dynamical systems, this invariant distribution is typically singular almost everywhere on the attractor, and it is necessary to introduce Gaussian noise into the system to make $\rho$ sufficiently smooth to apply the GFDT Gritsun and Branstator (2007) effectively.
@@
where $\bm{C}(\mathcal{T})$ is the correlation matrix, with elements $C_{ij}(\mathcal{T})=\langle u_{i}(\mathcal{T})u_{j}(0)\rangle_{\mathcal{J}}$. The property that $R_{ij}(0)=\delta_{ij}$, where $\delta_{ij}$ is the Kronecker delta, is a general property that holds for the discrete response matrix, see the Supplementary Information (SI). Constructing the correlation matrix is straightforward, allowing Equation (5) to serve as a convenient method for estimating the response matrix when the underlying steady-state distribution is well-approximated by a Gaussian distribution.
@@
Here, we employ score-based generative modeling to derive the response function directly from data. The approach uses denoising score matching to facilitate the parameterization of the dynamical system’s attractor score function directly, following methodologies suggested by Song et al. (Song and Ermon, 2019; Song et al., 2020). Specifically, we use a U-Net to represent the score-function Ronneberger et al. (2015); Bischoff and Deck (2023), but other choices of architecture are necessary depending on the application. Our approach only requires data and does not require approximating the steady-state distribution as Gaussian or additional numerical simulations for different perturbations. Refer to the SI for additional details on the architecture, training the model, obtaining the response function, and practical considerations.
Figure 1: Column 1: Example snapshots of the Three Systems. In the top row, Ornstein-Uhlenbeck PDE, in the middle row Allen-Cahn, in the bottom row Navier-Stokes. Columns 2-5: Response functions for Three Systems. The response to a perturbation at the pixel component $i=1$ is evaluated at various pixel components with coordinates indicated at each panel. All evaluated components align with the advection direction. Responses are computed by integrating the dynamical system (orange lines, “Dynamical”) via Gaussian approximation (black lines, “Gaussian”), using the score function derived from generative modeling (red lines, “Generative”), and using the analytic score function (blue lines, “Analytic”). Columns 6-7: Response Function Errors for the Three Systems. Here, we show both the RMS error and the Infinity Norm error of the Response function as a function of time. Here we use the Dynamic Response function as the “ground truth”. We see that the Generative response (red) generally outperforms the Gaussian response (blue) for the nonlinear cases (Allen-Cahn and Navier-Stokes), especially in the Infinity Norm, and performs similarly for the linear system (Ornstein-Uhlenbeck). The response function computed with the exact score does not have zero error relative to the dynamical response function because of errors introduced in numerical discretization and empirical averaging.
We apply the proposed method to three different dynamical systems. We compare the response function computed in this way (Generative) to a response computed by solving the perturbed dynamical system (Dynamical), a response using an analytically derived score function (Analytical) for the systems for which it is possible to derive one, and the response obtained through the Gaussian approximation (Gaussian).
@@
An Ornstein-Uhlenbeck Stochastic PDE. – Our first system is an Ornstein-Uhlenbeck (OU) system Uhlenbeck and Ornstein (1930) with spatially correlated noise, Maller et al. (2009); Griffin and Steel (2006); Frank et al. (2000); Vatiwutipong and Phewchean (2019). The steady-state trajectories of the system exhibit Gaussian statistics, which makes the Gaussian approximation exact, and hence serves as a baseline for testing our method.
@@
A modified Allen-Cahn system. – Our second system is the Allen-Cahn system, an example of a reaction-diffusion equation used for modeling phase separation in multi-component systems (see Allen and Cahn, 1972, who studied applications to binary alloys). Reaction-diffusion equations (Kolmogorov et al., 1937) are partial differential equations combining a diffusive term with a (non)linear reaction term. They are commonly used to study pattern formation (e.g., (Turing, 1990; Callahan and Knobloch, 1999; Kondo and Miura, 2010)) and chemical processes.
@@
For both the OU and Allen-Cahn systems, the analytic score function and the machine-learned score will differ by a constant factor due to the difference between the Kronecker delta and the Dirac delta functions. In both examples, the analytic score function is $32^{2}$ larger. We divide the analytic score by this factor for comparison with the machine-learned score. Refer to the SI for the derivation of the analytic score functions.
Navier-Stokes equation. – We next study a version of the 2D Navier-Stokes model as described in Flierl and Souza (2024), where we take the state variable to be the vorticity, $u$. This system was chosen as a stepping-stone towards the more complex fluids observed in climate systems; however, the vorticity statistics are not strongly non-Gaussian due to the chosen resolution as would be expected from a coarse-grained system, Sardeshmukh and Sura (2009). Hence, we expect good performance from the Gaussian approximation to the response function. The data-generating equation is
@@
$\mathcal{D}u$ $=-\nu_{h}\Delta^{-2}u-\nu\Delta^{2}u-\nu_{0}\int u\text{d}x\text{d}y$ (13)
@@
where $\mathcal{D}$ is a dissipation operator with constants ($\nu_{h}=10^{-2}$, $\nu=10^{-5}$, $\nu_{0}=(2\pi)^{-2}$), $A=\pi$ is a mean advection term, and $\varsigma$ is a random wave forcing from Flierl and Souza (2024). The random wave forcing introduces hidden dynamics into the system, thus the score function is estimated using only partial state information. The inverse Laplacian is defined to preserve the mean of the original variable.
@@
Results. – We first generate time series data corresponding to Equations (6),  (9), and Equation (11). These data are used to compute response functions and the training sets for the score model. To create the training data, we determine autocorrelation times for each and subsample the trajectories to create statistically independent snapshots. Figure (1) Column (1) shows training data samples of the numerically simulated systems. All details regarding the numerical simulations (time-stepping, spatial discretization), the datasets and the training of the score models are provided in the SI, along with an in-depth comparison between data samples generated via numerical simulation of the Allen-Cahn system and those produced using the learned score function of this system through reverse diffusion. Training the neural network took a few hours to a few days to for each score model on a single Nvidia A100.
@@
Owing to the periodic boundary conditions and symmetries of the systems under study, the response matrix elements exhibit translational invariance in the $x$ and $y$-directions. However, due to the advection term, the response function in the $x$-direction depends on both the distance and the direction (sign) between pixels while the $y$-direction only depends on the distance. This invariance implies that the index $i$ in the response matrix $R_{ij}$ can be arbitrarily fixed. We chose $j=1$ and calculated the response matrix elements along the direction of advection (denoted by $i$), as this is the direction information propagation, where we observe a more significant response to perturbations at pixel component $i=1$. The resulting responses of nearby pixels as a function of time for the three systems are presented in Figure (1) Columns (2-5). The generative model is consistent with the dynamical response function (“our ground truth”) in all cases. For the OU system, the Gaussian response function also matches the ground truth; this is expected as this system exhibits Gaussian statistics. For the Allen-Cahn system, it is clear that the Gaussian approximation is a poor approximation to the response function. And lastly, for the Navier-Stokes system we see improvement in using the generative score over the Gaussian response, despite the coarse Navier-Stokes system being quasi-Gaussian.
@@
In Figure (1) Columns (6-7), we compared the error over the entire state field between the Generative, Gaussian, and Analytic (for the systems where it is available) response functions with the Dynamical response function considered as “ground truth” for the three systems as a function of time. We used two different metrics to determine the errors: the $L^{2}$ and the $L^{\infty}$ norm. The $L^{2}$ norm measures the average error across all pixels in the domain, providing a sense of the overall error distribution. It is helpful in understanding the general accuracy of the response function over the entire field. On the other hand, the $L^{\infty}$ norm measures the maximum error at any single pixel, highlighting the worst-case error. This norm is particularly relevant to ensure the minimization of the largest deviations. From the comparison in Figure (1), it is evident that the Generative response function generally outperforms the Gaussian response in the nonlinear cases (Allen-Cahn and Navier-Stokes), especially in terms of the $L^{\infty}$ norm at early to intermediate times. This observation indicates that the Generative model is better at capturing the extreme responses at individual pixels. The Gaussian approximation is exact for the Ornstein-Uhlenbeck system, which has a Gaussian steady-state distribution, and both the Generative and Gaussian response functions perform similarly.
@@
Drawbacks and Practical Considerations. – While score-based generative models offer a powerful tool for estimating response functions, there are practical challenges. One issue is that training these models requires substantial data or a sufficiently good inductive bias in the parameterized model for the score function. Additionally, setting up and training neural networks can often require significant resources, engineering, and time. Expertise in training neural networks is essential. For example, suppose the data or training time is insufficient, and one is using a U-Net to represent the score function. In that case, the generative response function tends to match the Gaussian approximation’s performance due to the U-Net’s linear bypass connections. A further drawback with any machine-learning method is the lack of performance “out-of-sample”; indeed, in the SI, we demonstrate that the score model for the Allen-Cahn system performs poorly when evaluated on data dissimilar from the training data.
@@
For many applications, especially those mentioned in the introduction, several additional steps may need to be taken to apply the methods from this study effectively. One crucial step is the incorporation of domain-specific knowledge into the design of the neural network architecture. This will significantly improve the model’s inductive bias and alleviate the data burden. Additionally, domain-specific challenges such as non-stationarity or incomplete state information must be incorporated (Majda et al., 2010). This is particularly salient for climate data, which exhibits diurnal, seasonal, and longer timescale variations (Gershgorin and Majda, 2010; Majda and Qi, 2019). As always, it is crucial to validate the trained model with known benchmarks or analytical solutions.
@@
Conclusions. – In this study, we use score-based diffusion models to compute response functions, offering a purely data-driven method to calculate accurate response functions. We validated this approach on a system with Gaussian statistics (Ornstein-Uhlenbeck) and two dynamical systems with more complex distributions: a reaction-diffusion equation (Allen-Cahn) and the Navier-Stokes equations. The former is a class of SPDEs that have applications across biology, chemistry, and physics, and the latter is a system exhibiting some of the characteristics of more complex fluids like the Earth’s atmosphere or ocean. We compared the response function of these systems computed in multiple ways: using an analytic expression for the score, using the generative score function, using a dynamical approach by introducing perturbations to initial conditions and solving the equations of motion, and using a Gaussian approximation. For the nonlinear cases, we found that the Gaussian response function over-predicted the response to neighboring perturbations and that the generative response function agreed well with both the dynamical and analytical response functions. Compared with the Gaussian response, the RMSE of the generative response function had significantly reduced error. This suggests that score-based models are a useful tool for studying system response functions, especially when the dynamics are strongly nonlinear.